# The impact of data set similarity and diversity on transfer learning success in time series forecasting

### v1

@@ 4.3 Uncertainty estimation
Results resemble the accuracy results listed in Table 4: within the pre-trained models, the Multisource data set represents the best source data set for transfer learning. As for the accuracy, the uncertainty estimation usually improves with fine-tuning, in 74% of the cases. While the zero-shot performance of the Electricity source model is inferior to other pre-trained models, it catches up with fine-tuning. The Traffic data set is especially hard to forecast.
To summarize: comparing the pre-trained models, those pre-trained on the Multisource data set provides superior uncertainty estimation performance and fine-tuning enhances the performance in general.
4.4 Training time
Source / Target Solar Traffic Kaggle Wholesaler2 Wholesaler1 Sum
@@ 4.4 Training time
5 Interrelations between source and target data sets
The overall goal of this paper is to find relations between source-target similarity, source diversity, and transfer learning performance in accuracy, bias and uncertainty estimation. Having evaluated the forecasting performances for zero-shot and fine-tuned pre-trained models in Section 3, we turn to measuring data set similarity and diversity in this section in order to map these characteristics of the source and target data sets to the performance of transfer learning in the next section.
To that end, we compute two different time series feature sets as well as representative time series for all source and target data sets. Based on these, we calculate two feature-based and one shape-based similarity measure, and two feature-based diversity measures. As feature-based similarity measure, we use the median Euclidean feature distances between all sources and targets. We express the feature-based diversity per data set by the feature variance and visualize this in the latent principal component space. To not only concentrate on feature-based similarity, we also include Dynamic Time Warping (DTW) distances of the data set representatives. In the following, we describe the time series features, similarity and diversity measures in more detail.
We restrict the number of time series considered per data set for these calculations by randomly drawing their ids. If a data set contains less than this number of time series, all are taken into account. We limit the number of time series to save computational resources while still producing representative results.
5.1 Time series features
We extract ten common time series features to characterize the data via the tsfresh package (Christ et al.,, 2018): ”absolute energy”, ”intermittency”, ”mean”, ”median”, ”kurtosis”, ”skewness”, ”standard deviation”, ”aggregated autocorrelation”, ”lumpyness” and ”aggregated linear trend”. This selection is parsimonious and covers both scale and shape characteristics. To not limit our analysies to a single hand-selected feature set, and since catch22 features (Lubba et al.,, 2019) are known to be a helpful input for time series classification, we further extract catch22 features via the pycatch22 toolkit  [^1] https://github.com/DynamicsAndNeuralSystems/pycatch22. Further details on the extracted features can be found in the appendix.
5.2 Feature-based data set similarity
To assess the similarity of source and target data sets, we z-transform the extracted features for each data set, so that they have zero mean and unit variance. To save computational effort, we only compute these transformed features for at most 100 randomly selected time series per data set. Based on that, we calculate the Euclidean distance between each source and target time series per feature and, subsequently, compute the median over all time series distances per feature. Finally, we compute the median over the feature distances for every source-target data set pair. The resulting distance matrix provides the median feature distance between each source and target data set. We compute these distance matrices for both the tsfresh and the catch22 features.
Source/Target Solar Traffic Kaggle Wholesaler2 Wholesaler1
Multisource 1.23 0.31 0.29 0.64 0.38
M5 1.16 0.31 0.29 0.62 0.37
M4 1.20 0.38 0.38 0.71 0.50
Electricity 0.87 0.47 0.56 0.79 0.71
NN5 1.36 0.29 0.28 0.53 0.31
Exchange Rate 0.95 0.30 0.28 0.67 0.31
Source/Target Solar Traffic Kaggle Wholesaler2 Wholesaler1
Multisource 0.76 0.73 0.68 0.68 0.66
M5 0.81 0.73 0.65 0.63 0.64
M4 1.15 0.96 1.01 1.23 1.05
Electricity 0.55 0.52 0.61 0.95 0.88
NN5 0.72 0.67 0.76 0.66 0.66
Exchange Rate 1.15 1.08 0.92 1.17 1.08
Table 8 lists the median Euclidean feature distances between the source and target data sets for the tsfresh features, while Table 9 lists those for catch22. Smaller values indicate more similarity. In terms of tsfresh feature distances, NN5 is the closest and Electricity is the most distant to all target data sets except Solar Energy, where it is the other way round. In terms of catch22 features, Electricity is the closest source data set to all target, except for the wholesalers, where it is the M5. This shows that the feature sets calculate fundamentally different time series characteristics. Hence, we keep them both for further analysis.
5.3 Shape-based data set similarity
To also account for shape-based similarity measures, we compute a representative time series of each source and target data set based on the DTW Bary Center Averaging (DBA) method (Petitjean et al.,, 2011). In that, we use the tslearn package (Tavenard et al.,, 2020), which was implemented based on Schultz and Jain, (2018). For this step the time series were first scaled to zero mean and unit variance per data set and then at most 1,000 time series were randomly drawn per data set.
Source/Target Solar Traffic Kaggle Wholesaler2 Wholesaler1
Multisource 0.69 2.48 0.19 3.31 3.29
M5 5.35 4.23 6.03 4.03 3.78
M4 27.01 26.13 28.79 9.51 8.96
Electricity 4.02 2.20 4.55 3.31 3.09
NN5 3.52 3.30 3.82 3.44 3.36
Exchange Rate 7.84 7.37 8.62 3.07 3.80
The distances between the data sets are then assessed by the DWT (Petitjean et al.,, 2011) distances of the data set representatives, implemented by DTAIDistance (Meert et al.,, 2020). They are provided in Table 10. Here, either the Electricity or the Multisource are the most similar source for all target data sets, except for Wholesaler2 where the Exchange Rate data set is the closest.
5.4 Source data set diversity
To measure the diversity per source data set, we consider the variance of the two feature sets calculated in Section 5.1 for at most 1,000 time series per data set. The variance of each feature is Min-Max scaled between zero and one over the source data sets. Hence, the source with the highest variance of a feature in comparison to the other sources, is indicated by a one for this feature. To compress the information, the sum over the feature variances is built and saved as diversity value for the respective source data set.
Source tsfresh catch22
Multisource 1.85 8.89
M5 1.28 7.68
M4 4.01 12.71
Electricity 5.44 5.78
NN5 0.54 4.68
Exchange Rate 0.03 6.30
5.5 Visualization of data set diversity
To visualize the data set diversity, we implement a principal component analysis (PCA) of the time series features of both sets. We randomly select 1,000 time series and standardize their extracted features over all data sets, so that they get projected into the same latent space. We compute the PCA with two principal components via the sklearn package (Pedregosa et al.,, 2011).
(a) Sources Tsfresh
(b) Targets Tsfresh
(c) Sources Catch22
(d) Targets Catch22
The bottom left image of Figure 2 shows the source data sets according to their catch22 features in the PCA space. The picture is completely different to that of the tsfresh features, as are the respective source-diversity values in Table 11. It demonstrates that the M4 data set is quite different from the others in terms of location and spread, which is in line with being the most diverse according to Table 11. The bottom right image of Figure 2 visualizes the target data sets according to the catch22 features together with the Multisource. As for the tsfresh features, the wholesales data sets occupy the same space. Again, the time series of all sources, represented by the Multisource, do not cover all the space of the targets.
Given the analysis of similarities between source and target data sets and of diversities of the sources, we now assess their relations to transfer learning forecasting performance on the targets. To that end, we create separate plots with each showing one link between transfer learning accuracy (see Table 4), bias (see Table 5), uncertainty estimation (see Table 6) and similarity by tsfresh feature distance (see Table 8), catch22 feature distance (see Table 9), DTW distance (see Table 10), source diversity by tsfresh, source diversity by catch22 features (for both see Table 11), for zero-shot and fine-tuning forecasts for all target data sets.
The plots further discussed in the following indicate a roughly linear relationship throughout the targets. They only include zero-shot forecasting performance, since fine-tuning harmonizes forecasting performance, as seen in Section 4, which consequently flattens the trends in their relation to similarity and diversity values. This suggests that fine-tuning makes forecasts more independent of the source data set used.
6.1 The effect of source-target similarity
Figure 3 shows that source data sets that are more similar to the target data sets in their tsfresh features create more accurate forecasts and less forecast bias. This can be seen by the target curves going from lower forecast error and bias to higher ones as the distance to the sources increases. The relation to the forecast error is more distinct for the Traffic and Wholesaler1 dataset than for the other targets. The Kaggle data set constitutes an exception with regard to the forecast bias. Also, we state the relations in a relative manner, as the Solar data set for example has the largest distance to the source data sets but does not provide the worst forecast performance. We could not find such consistent relations regarding the uncertainty estimation, catch22 features or DTW distances.
(a) Tsfresh similarity and zero-shot forecast error
(b) Tsfresh similarity and zero-shot forecast bias
Figure 3: Relations between tsfresh source-target distance and zero-shot transfer learning accuracy and bias. Each curve corresponds to a target data set and each point to a source data set with their respective tsfresh feature distance on the x-axis and the zero-shot transfer learning accuracy or bias on the y-axis.
6.2 The effect of source diversity
Figure 4(a) and (c) demonstrate that more diversity in catch22 features is favorable for zero-shot accuracy and uncertainty estimation. This can be seen by the target curves going from higher forecast errors and uncertainty estimation errors to lower ones with increasing source diversity. As for the source-target similarity results, Traffic and Wholesaler1 provide a broader range of error values. The two most right points of each curve correspond to the Multisource and M4 source data sets, according to Table 11, providing superior accuracy and bias performance in general. Further, subfigure (b) shows that sources with a higher diversity in tsfresh features produce transfer learning forecasts with higher bias. A look at the underlying Tables 11 and 5 shows that e.g. the Electricity source data set, which has the highest tsfresh diversity, shows the highest zero-shot transfer learning forecast biases over the targets. The Solar and Kaggle data sets show more of a flat progression, but do not show the opposite of the other curves, and therefore are an exception. As the diversity measure is defined in relation to the regarded source data sets, the results also need to be interpreted relatively.
(a) Catch22 diversity and zero-shot forecast error
(b) Tsfresh diversity and zero-shot bias
(c) Catch22 diversity and zero-shot uncertainty estimation
Figure 4: Relation between the catch22 (a+c) and tsfresh (b) source diversity and their zero-shot transfer learning forecast error (a), bias (b) and uncertainty estimation error (c). Each curve corresponds to a target data set and each point to a source data set with its catch22 diversity on the x-axis and their zero-shot transfer learning forecasting performance on the y-axis.
7 Conclusion
7.1 Summary
The goal of this study was to find characteristics of source and target data sets to be beneficial for transfer learning forecasting performance. Therefore, zero-shot and fine-tuned forecasting performance regarding accuracy, bias and uncertainty estimation of five pre-trained single-source models, their ensemble and the foundation model trained on the combination of the source data sets, on five diverse real-world target data sets was assessed. After that, the relations of source and target data sets were analyzed by calculating two feature-based and one shape-based similarity measures. The diversity of the source data sets was computed as the variance of the two feature sets and visualized by projecting the data sets into the latent PCA space.
The following relations of source-target similarity and source diversity to the forecasting performance were found by plotting them against each other: A high similarity between the source and target data set in tsfresh features enhances the zero-shot accuracy and bias. A high diversity of the source data set in catch22 features leads to better zero-shot accuracy and uncertainty estimation, while high diversity in tsfresh features increases zero-shot bias. These relationships are much weaker when the pre-trained model is fine-tuned on the target data, since fine-tuning harmonizes the forecasting performance.
Summarizing the results of the analyses documented here, we can state a distinct set of insights:
  • 1.
    Fine-tuning smoothens the relations between similarity, diversity and forecasting performance. Hence, it is more relevant to investigate data set characteristics in regard to transfer learning success in the zero-shot case. If there are resources, we recommend to fine-tune pre-trained models for enhanced transfer learning accuracy and uncertainty estimation; then the source data set does not play a major role, according to our findings.
  • 2.
    The feature-based similarity and diversity measures reveal different relations to the forecasting performance and are therefore both helpful in choosing a source data set out of a variety. According to our findings, we recommend choosing the source data set which is either the most similar in tsfresh features or the most diverse in catch22 features. For the latter, our pre-trained model on the M4 data can be used.
  • 3.
    The DTW distance measure did not show any relation to the forecasting results.
  • 4.
    Our results justify the approaches in the literature of selecting similar source data sets for a specific target or building a foundation model with diverse source data sets.
  • 5.
    Regardless of the source domain, similarities in tsfresh features can be an indicator of zero-shot transfer learning success.
7.2 Limitations and research directions
The source data sets considered in our study study were not chosen to be similar to any of the target data sets or to be as diverse as possible. Hence, our results rely on random source data sets, of which we can measure their similarity to the targets and their diversity only in a relative manner, while the closest source in our experiment might still be not similar in an absolute sense and the most diverse source may not be diverse in comparison to much larger data sets used for training foundation models. Therefore, we can only supply relative results and no similarity or diversity metric with absolute thresholds telling whether a source should be used or not to guarantee for transfer learning success.
Further, the models and results presented here are restricted to weekly time series data. Including multiple granularities and also transfer learning across them was outside the scope of this paper. We also excluded multivariate time series or external regressors, since the goal was to identify relations between similarities and diversities to forecasting performance and not relations between time series. Nevertheless, it would be interesting to see how the additional information, which is contained in a couple of data sets of the Monash Forecasting Repository (Godahewa et al.,, 2021), changes the relations. Furthermore, we used the same forecast horizon of fifteen weeks for all data sets for training and testing. Nevertheless, the pre-trained DeepAR models applied and published (upon acceptance) for transfer learning are able to do zero-shot and fine-tuned forecasts also for other horizons. The results might be different for other horizons. We leave these topics for further research.
Acknowledgements
Special thanks go to Benedikt Sonnleitner, Leif Fedderson and Ali Ismail-Fawaz for thoughtful discussions on the ideas around this paper. Also we would like to thank Julia Schemm, Nico Beck and Vrinda Gupta who co-setup the basic code on which the experiments were built upon. We extend our sincere gratitude to the providers of the open datasets, whose generous sharing of data resources has been instrumental in the advancement of our research.
Appendix
Name
Explanation
Source
Absolute Energy
Sum of squared values
tsfresh
Intermittency
Length of time series divided
@@ Acknowledgements
Mean
Mean of the time series
tsfresh
Median
Median of the time series
tsfresh
Kurtotsis
Kurtosis of the time series
tsfresh
Skewness
Skewness of the time series
tsfresh
Standard deviation
Standard deviation of the time series
tsfresh
Aggregate autocorrelation max
Maximal autocorrelation of the first five lags of the time series
tsfresh
Lumpyness
Variance of non-zero time steps divided
@@
Linear Trend
Slope of the time series
tsfresh

### v2

@@ 4.3 Uncertainty estimation
The results resemble the accuracy results for point forecasts in Table 4: within the pre-trained models, the Multisource data set represents the best source data set for transfer learning. Similar to the accuracy, the uncertainty estimation usually improves with fine-tuning, in 74% of the cases. While the zero-shot performance of the Electricity source model is inferior to other pre-trained models, it catches up with fine-tuning. To summarize: comparing the pre-trained models, those pre-trained on the Multisource data set provides superior uncertainty estimation performance and fine-tuning enhances the performance in general.
5 Interrelations between source and target data sets
The overall goal of this paper is to find relations between source-target similarity, source diversity, and transfer learning performance in accuracy, bias and uncertainty estimation. Having evaluated the forecasting performances for zero-shot and fine-tuned pre-trained models in Section 4, we turn to measuring data set similarity and diversity. This then allows us to map these characteristics of the source and target data sets to the performance of transfer learning in the next section.
We compute two different time series feature sets as well as averaged time series for all source and target data sets. Based on these, we calculate two feature-based and one shape-based similarity measure, and two feature-based diversity measures. As feature-based similarity measure, we use the median Euclidean feature distances between all sources and targets. We express the feature-based diversity per data set by the feature variance and visualize this in the latent principal component space. We also include DTW distances of the data set representatives. In the following, we describe the time series features, similarity and diversity measures in more detail.
We restrict the number of time series considered per data set for these calculations by randomly drawing 1,000 of their ids. If a data set contains less than this number of time series, all are taken into account. We limit the number of time series to save computational resources while still producing representative results.
5.1 Time series features
We extract ten common time series features to characterize the data via the tsfresh package (Christ et al., 2018): ”absolute energy”, ”intermittency”, ”mean”, ”median”, ”kurtosis”, ”skewness”, ”standard deviation”, ”aggregated autocorrelation”, ”lumpyness” and ”aggregated linear trend”. This selection is parsimonious and covers both scale and shape characteristics. Further details on the extracted features can be found in the Appendix in Table 14. To not limit our analysis to a single hand-selected feature set, and since catch22 features (Lubba et al., 2019) are known to be a helpful input for time series classification, we further extract catch22 features via the pycatch22 toolkit  [^1] https://github.com/DynamicsAndNeuralSystems/pycatch22. These include e.g. key figures on the distribution, temporal statistics and linear and non-linear autocorrelation. We use all 22 catch22 features plus the ”mean” and ”standard deviation”.
5.2 Feature-based data set similarity
To assess the similarity of source and target data sets, we z-transform the extracted features over all source data sets and separately over all target data sets, so that they have zero mean and unit variance per feature. Let $f_{i,s_{n}}$ be the standardized feature vector of the $i$-th time series of the source data set $s_{n}$ and $f_{j,t_{m}}$ the standardized feature vector of time series j of the target data set $t_{m}$, either for tsfresh or catch22. Based on that, we calculate the Euclidean distance between each feature $p$ of each source and target time series pair, $dist(f_{i,s_{n}}[p],f_{j,t_{m}}[p])$. Then, we compute the median over all time series pairs’ distances per feature $p$: $dist_{p}(s_{n},t_{m})=median(dist(f_{i,s_{n}}[p],f_{j,t_{m}}[p]))$. Finally, we compute the median over the feature distances for every source-target data set pair: $dist(s_{n},t_{m})=median(dist_{p}(s_{n},t_{m}))$. The resulting distance matrix $dist(s_{n},t_{m})$ provides the median feature distance between each source and target data set. We compute these distance matrices for both the tsfresh and the catch22 features.
Source/Target Solar Traffic Kaggle Wiki Wholesaler2 Wholesaler1
Multisource 1.36 0.24 0.22 0.46 0.30
M5 1.35 0.24 0.22 0.45 0.29
M4 1.50 0.26 0.25 0.59 0.41
Electricity 1.22 0.30 0.29 0.50 0.31
NN5 1.64 0.21 0.20 0.42 0.25
Exchange Rate 1.23 0.19 0.18 0.46 0.25
Source/Target Solar Traffic Kaggle Wiki Wholesaler2 Wholesaler1
Multisource 0.82 0.73 0.65 0.63 0.67
M5 0.82 0.71 0.64 0.61 0.65
M4 1.22 1.03 1.07 1.25 1.1
Electricity 0.59 0.61 0.65 0.97 0.91
NN5 0.75 0.71 0.74 0.73 0.69
Exchange Rate 1.29 1.09 1.05 1.22 1.17
Table 7 lists the median Euclidean feature distances between the source and target data sets for the tsfresh features, and Table 8 lists those for catch22. Smaller values indicate higher similarity. In terms of tsfresh feature distances, NN5 or Exchange Rate is the closest to all target data sets and Electricity or M4 is the most distant except Solar Energy, where it is the other way round. In terms of catch22 features, Electricity is the closest source data set to Solar and Traffic and otherwise it is the M5. This highlights that the feature sets capture fundamentally different time series characteristics. Hence, we keep them both for further analysis.
5.3 Shape-based data set similarity
To also account for shape-based similarity measures, we compute a representative time series of each source and target data set based on the DTW Bary Center Averaging (DBA) method (Petitjean et al., 2011). In that, we use the tslearn package (Tavenard et al., 2020), which was implemented based on Schultz and Jain (2018). The time series are first scaled to zero mean and unit variance per data set.
Source/Target Solar Traffic Kaggle Wiki Wholesaler2 Wholesaler1
Multisource 0.69 2.48 0.19 3.31 3.29
M5 5.35 4.23 6.03 4.03 3.78
M4 27.01 26.13 28.79 9.51 8.96
Electricity 4.02 2.20 4.55 3.31 3.09
NN5 3.52 3.30 3.82 3.44 3.36
Exchange Rate 7.84 7.37 8.62 3.07 3.80
The distances between the data sets are then assessed by the DTW (Petitjean et al., 2011) distances of the data set representatives, implemented by DTAIDistance (Meert et al., 2020). They are provided in Table 9. Here, either the Electricity or the Multisource are the most similar source for all target data sets, except for Wholesaler2 where the Exchange Rate data set is the closest.
5.4 Source data set diversity
To measure the diversity per source data set, we consider the variance of the two feature sets calculated in Section 5.1. The variance of each feature is Min-Max scaled between zero and one over the source data sets. Hence, the source with the highest variance of a feature in comparison to the other sources, is indicated by a one for this feature. Then the sum over the feature variances is built and saved as diversity value for the respective source data set.
Source tsfresh catch22
Multisource 1.73 8.95
M5 1.54 8.25
M4 4.01 12.86
Electricity 5.51 6.08
NN5 0.69 5.15
Exchange Rate 0.03 6.34
5.5 Visualization of data set diversity
To visualize the data set diversity, we implement a two-dimensional principal component analysis (PCA) of the time series features of both sets. We standardize their features over all data sets, so that they get projected into the same latent space. We compute the PCA with two principal components via the sklearn package (Pedregosa et al., 2011).
(a) Sources Tsfresh
(b) Targets Tsfresh
(c) Sources Catch22
(d) Targets Catch22
The bottom left image of Figure 3 shows the source data sets according to their catch22 features in the PCA space. The picture is completely different to that of the tsfresh features, as are the respective source diversity values in Table 10. It demonstrates that the M4 data set is quite different from the others in terms of location and spread, which is in line with being the most diverse according to Table 10. The bottom right image of Figure 3 visualizes the target data sets according to the catch22 features together with the Multisource. As for the tsfresh features, the wholesales data sets occupy the same space. Again, the time series of all sources, represented by the Multisource, do not cover all the space of the targets.
Given the analysis of similarities between source and target data sets and of diversities of the sources, we now assess their relations to transfer learning forecasting performance on the targets. To that end, we create separate plots with each showing one link between one of the transfer learning accuracy (see Table 4), bias (see Table 5) and uncertainty estimation (see Table 6) measures and one of the similarity by tsfresh feature distance (see Table 7), catch22 feature distance (see Table 8), DTW distance (see Table 9), source diversity by tsfresh, source diversity by catch22 features (for both see Table 10) measures, for either zero-shot or fine-tuning forecasts for all target data sets. To test for linear relationships, we fit a linear regression to each target relation line and report the slope and its p-value.
In the following, we show only those plots together with their linear regression fitting information for which we find a roughly linear relationship throughout the targets. They only include zero-shot forecasting performance, since fine-tuning harmonizes forecasting performance, as seen in Section 4, which consequently flattens the trends in their relation to similarity and diversity values. This suggests that fine-tuning makes forecasts more independent of the source data set used.
6.1 The effect of source diversity
(a) Catch22 diversity and zero-shot forecast accuracy error
(b) Tsfresh diversity and zero-shot bias
(c) Catch22 diversity and zero-shot uncertainty estimation
Figure 4: Relation between the catch22 (a+c) and tsfresh (b) source diversity and their zero-shot transfer learning forecast accuracy error (a), bias (b) and uncertainty estimation error (c). Each curve corresponds to a target data set and each point to a source data set with its respective diversity on the x-axis and its zero-shot transfer learning forecasting performance on the y-axis.
Target Acc_slope P_val Unc_slope P_val Bias_slope P_val
Wholesaler1 -0.29 0.34 -7.88 0.20 0.31 0.01
Wholesaler2 -0.07 0.01 -0.28 0.21 0.41 0.00
Kaggle -0.04 0.36 -1.66 0.24 -1.66 0.24
Traffic -0.30 0.44 -5.95 0.71 -5.95 0.71
Solar -0.01 0.82 -0.34 0.70 -0.34 0.70
Table 11: Linear regression fit of catch22 diversity and zero-shot forecast accuracy and uncertainty error, and tsfresh diversity and zero-shot forecast bias, per target data set. Statistically significant slope values are printed in bold.
Figure 4(a) and (c) demonstrate that more diversity in catch22 features is favorable for zero-shot accuracy and uncertainty estimation. This can be seen by the target curves going from higher forecast accuracy errors and uncertainty estimation errors to lower ones with increasing source diversity. Traffic and Wholesaler1 provide a broader range of error values. The Solar and Kaggle Wiki data sets show more of a flat progression, but do not show the opposite of the other curves. When fitting a linear regression per target data set, all slopes are negative in the columns of Table 11, which correspond to accuracy and uncertainty, indicating that more diversity improves forecast accuracy and uncertainty estimation. Note that for the accuracy, only for Wholesaler2 the p-values of the slope are statistically significant. For the uncertainty none of the slope coefficients are significant.
Further, Subfigure (b) shows that sources with a higher diversity in tsfresh features produce transfer learning forecasts with higher bias. A look at the underlying Tables 10 and 5 shows that e.g. the Electricity source data set, which has the highest tsfresh diversity, shows the highest zero-shot transfer learning forecast biases over the targets. The linear regression curves show almost always a positive slope (bias columns in Table 11), which are significant for the wholesalers.
6.2 The effect of source-target similarity
Figure 5: Relations between tsfresh source-target distance and zero-shot transfer learning bias. Each curve corresponds to a target data set and each point to a source data set with its respective tsfresh feature distance on the x-axis and the zero-shot transfer learning bias on the y-axis.
Target Bias_slope P_val
Wholesaler1 7.90 0.14
Wholesaler2 10.69 0.11
Kaggle 0.04 1.00
Traffic 34.04 0.06
Solar 0.57 0.84
Table 12: Linear regression fit of tsfresh distance and zero-shot forecast bias per target data set.
Figure 5 shows that source data sets that are more similar to the target data sets in their tsfresh features create less forecast bias. This can be seen by the target curves going from lower forecast bias to higher ones as the distance to the sources increases. All linear regression curves provide a positive slope (Table 12), which is not statistically significant.
Note that the relations are relative for the source data sets of each target, as we scale over the source data sets to get similarity and diversity measures. Hence, the relations are not absolute, which can be seen at the Solar data set for example, which has the largest distance to the source data sets but does not provide the worst forecast performance.
We could not find such consistent relations regarding the accuracy, uncertainty estimation, catch22 features or DTW distances.
7 Conclusion
To conclude, we provide a summary of the methods and results, as well as limitations and directions for further research.
7.1 Summary
We investigated the impact of source-target data set similarity and source data set diversity onto the success of transfer learning. Therefore, we assessed zero-shot and fine-tuned forecasting performance regarding accuracy, bias and uncertainty estimation of five pre-trained single-source models, their ensemble and the foundation model trained on the combination of the source data sets, on five diverse real-world target data sets. After that, we analyzed the relations of source and target data sets by calculating two feature-based and one shape-based similarity measures. The diversity of the source data sets was computed as the variance of the two feature sets and visualized by projecting the data sets into the latent PCA space.
We find that high similarity between the source and target data set in tsfresh features enhances the zero-shot bias. A high diversity of the source data set in catch22 features leads to better zero-shot accuracy and uncertainty estimation, while high diversity in tsfresh features increases zero-shot bias. As opposed to the feature-based similarity measures, the shape-based DTW distance measure does not show any relation to the forecasting results. Fine-tuning smoothens the relations between similarity, diversity and forecasting performance. Hence, it is more relevant to investigate data set characteristics in regard to transfer learning success in the zero-shot case. If there are resources, we recommend to fine-tune pre-trained models for enhanced transfer learning accuracy and uncertainty estimation; then the source data set does not play a major role, according to our findings. As stated above, feature-based similarity and diversity measures reveal different relations to the forecasting performance and are therefore both helpful in choosing a source data set out of a variety. According to our findings, we recommend choosing the source data set which is either the most similar in tsfresh features or the most diverse in catch22 features. For the latter, our pre-trained model on the M4 data can be used. Our results justify the approaches in the literature of selecting similar source data sets for a specific target or building a foundation model with diverse source data sets. Regardless of the source domain, similarities in tsfresh features can be an indicator of zero-shot transfer learning success.
7.2 Limitations and research directions
The source data sets considered in our study were not chosen to be similar to any of the target data sets or to be as diverse as possible. Hence, our results rely on random source data sets, of which we can measure their similarity to the targets and their diversity only in a relative manner, while the closest source in our experiment might still be not similar in an absolute sense and the most diverse source may not be diverse in comparison to much larger data sets used for training foundation models. Therefore, we can only supply relative results and no similarity or diversity metric with absolute thresholds telling whether a source should be used or not to guarantee for transfer learning success.
Many of the results shown in Section 6, especially the p-values of the regression lines, are not statistically significant. A reason might be the number of data points. Hence, we call for further studies examining these relationships with more source and target data sets.
Further, the models and results presented here are restricted to weekly time series data. Including multiple granularities and also transfer learning across them is outside the scope of this paper. We also exclude multivariate time series or external regressors, since the goal is to identify relations between similarities and diversities to forecasting performance and not relations between time series. Nevertheless, it would be interesting to see how the additional information, which is contained in a couple of data sets of the Monash Forecasting Repository (Godahewa et al., 2021), changes the relations. Furthermore, we use the same forecast horizon of fifteen weeks for all data sets for training and testing. Nevertheless, the pre-trained DeepAR models applied and published (upon acceptance) for transfer learning are able to do zero-shot and fine-tuned forecasts also for other horizons. The results might be different for other horizons. We leave these topics for further research.
Acknowledgements
Special thanks go to Leif Feddersen, Ali Ismail-Fawaz and Julius Mehringer for thoughtful discussions on the ideas around this paper. Also we would like to thank Julia Schemm, Nico Beck and Vrinda Gupta who co-setup the basic code on which the experiments were built upon. We extend our sincere gratitude to the providers of the open datasets, whose generous sharing of data resources has been instrumental in the advancement of our research.
Appendix
Training time
Source / Target Solar Traffic Kaggle Wiki Wholesaler2 Wholesaler1 Sum
@@ Acknowledgements
Tables
Name Explanation Source
Absolute Energy Sum of squared values tsfresh
Intermittency Length of time series divided
@@
Mean Mean of the time series tsfresh
Median Median of the time series tsfresh
Kurtotsis Kurtosis of the time series tsfresh
Skewness Skewness of the time series tsfresh
Standard deviation Standard deviation of the time series tsfresh
Aggregate autocorrelation max Maximal autocorrelation of the first five lags of the time series tsfresh
Lumpyness Variance of non-zero time steps divided
@@
Linear Trend Slope of the time series tsfresh